OpenStudy (anonymous):

Express y=3x^2-9x +1 in the form y=a(x+b)^2+c so i completed the square, but i got y=3(x-3/2)^2-77/4 .... its the wrong answer ><

OpenStudy (anonymous):

@ganeshie8 d'you know anything about completing squares for graphs? :3

ganeshie8 (ganeshie8):

\(\large y=3x^2-9x +1 \) \(\large y=3(x^2-3x) +1 \) \(\large y=3(x^2-3x + \frac{9}{4}) +1 - 3\times \frac{9}{4} \) \(\large y=3(x - \frac{3}{2})^2 +1 - \frac{27}{4} \) \(\large y=3(x - \frac{3}{2})^2 - \frac{23}{4} \)

ganeshie8 (ganeshie8):

your last term is wrong that all.. go through above to see where exactly you did the mistake

OpenStudy (anonymous):

hang on... why 3 x 9/4? o-o isn't it 1-(-9/2)^2?

ganeshie8 (ganeshie8):

inside the parenthesis, you're adding `9/4` right ?

OpenStudy (anonymous):

ohh shoot the square method can only be done if the coefficient of x^2 is one, right?

ganeshie8 (ganeshie8):

it still works here also notice that 9/4 = (3/2)^2

OpenStudy (anonymous):

yes, but i've never done a rule that needed me to multiply it by the x coefficient ><

ganeshie8 (ganeshie8):

inside the parenthesis you're adding (3/2)^2 so you need to subtract that outside the parenthesis : -(3/2)^2 however, since the whole parenthesis was multiplied by 3, you need to subtract : -3 (3/2)^2 = -3*9/4

OpenStudy (anonymous):

i only factorised 3 from the 3(x-3/2)^2, so shouldn't i only add it back to that one?

ganeshie8 (ganeshie8):

one way to make sense of above is to multiply again, simplify and see if you get back to the earlier equation

OpenStudy (anonymous):

okay.. so when i complete the square for an equation with the x^2 coefficient more than 1, i have to multiply the back number with the coefficient?

OpenStudy (anonymous):

what just happened o-o sorry i forgot to reply i fell asleep /shot

OpenStudy (anonymous):

wait and why the talk about code o-O

ganeshie8 (ganeshie8):

your thread was hijacked by @dan815 earlier

ganeshie8 (ganeshie8):

suppose you have : \(\large a(x+b)\)

ganeshie8 (ganeshie8):

if you add 2 and subtract inside the parenthesis, nothing changes : \(\large a(x+b) = a(x+b\color{red}{+2-2})\) right ?

ganeshie8 (ganeshie8):

by distributing, you can kick that -2 out of the parenthesis : \(\large a(x+b) = a(x+b\color{red}{+2}) \color{Red}{-} a(\color{red}{2})\)